# Pretraining Decision Transformers with Reward Prediction for In-Context Multi-task Structured Bandit Learning

### v2

@@ Abstract
In this paper, we study the multi-task structured bandit problem where the goal is to learn a near-optimal algorithm that minimizes cumulative regret. The tasks share a common structure and any optimal algorithm should exploit the shared structure to minimize the cumulative regret for an unseen but related test task. We use a transformer as a decision-making algorithm to learn this shared structure so as to generalize to the unseen test task. The prior work of pretrained decision transformers like DPT requires access to the optimal action during training which may be hard in several scenarios. Diverging from these works, our learning algorithm does not need the knowledge of optimal action per task during training but predicts a reward vector for each of the actions using only the observed offline data from the diverse training tasks. Finally, during inference time, it selects action using the reward predictions employing various exploration strategies in-context for an unseen test task. We show that our model outperforms other methods like DPT, and Algorithmic Distillation (AD) and matches the performance of algorithms that requires privileged information on the structure of the problem. Interestingly, we show that our algorithm, without the knowledge of the underlying problem structure, can learn a near-optimal policy in-context by leveraging the shared structure across diverse tasks. We show that when the shared structure breaks down with the introduction of new actions both during training and test time, our proposed algorithm fails to learn the underlying latent structure. We further show that our algorithm conducts an implicit two-phase exploration and validate all of these findings over several experiments spanning linear, non-linear, real-life datasets, bilinear, and latent bandit settings. Finally, we theoretically analyze the performance of our algorithm and obtain generalization bounds in the in-context multi-task learning setting.
@@ 1 Introduction
    We introduce a new pre-training procedure of learning the underlying reward structure and a decision algorithm. Moreover, PreDeToR by predicting the next reward for all arms circumvents the issue of requiring access to the optimal (or approximately optimal) action during training time.
@@ 2.1 Preliminaries
Define $[n]=\{1,2,\ldots,n\}$. In a multi-task structured bandit setting the expected reward for each action in each task is assumed to be an unknown function of the hidden parameter and action features (Lattimore & Szepesvári, 2020; Gupta et al., 2020). The interaction proceeds iteratively over $n$ rounds for each task $m\in[M]$. At each round $t\in[n]$ for each task $m\in[M]$, the learner selects an action $I_{m,t}\in\mathcal{A}$ and observes the reward $r_{m,t}=f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})+\eta_{m,t}$, where $\bm{\theta}_{m,*}\in\mathbb{R}^{d}$ is the hidden parameter specific to the task $m$ to be learned by the learner. The function $f(\cdot,\cdot)$ is the unknown reward structure. This can be $f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})=\mathbf{x}(I_{m,t})^{\top}\bm{\theta}_{m,*}$ for the linear setting or even more complex correlation between features and $\bm{\theta}_{m,*}$ (Filippi et al., 2010; Abbasi-Yadkori et al., 2011; Riquelme et al., 2018; Lattimore & Szepesvári, 2019; Dong et al., 2021).
@@ 2.2 In-Context Learning Model
Pretraining: Let $\mathcal{T}_{\text{pre}}$ denote the distribution over tasks $m$ at the time of pretraining. Let $\mathcal{D}_{\text{pre}}$ be the distribution over all possible interactions that the $\pi^{w}$ can generate. We first sample a task $m\sim\mathcal{T}_{\text{pre }}$ and then a context $\mathcal{H}_{m}$ which is a sequence of interactions for $n$ rounds conditioned on the task $m$ such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|m)$. So $\mathcal{H}_{m}=\{I_{m,t},r_{m,t}\}_{t=1}^{n}$. We call this dataset $\mathcal{H}_{m}$ an in-context dataset as it contains the contextual information about the task $m$. We denote the samples in $\mathcal{H}_{m}$ till round $t$ as $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This dataset $\mathcal{H}_{m}$ can be collected in several ways: (1) random interactions within $m$, (2) demonstrations from an expert, and (3) rollouts of an algorithm. Finally, we train a causal GPT-2 transformer model $\mathrm{TF}$ parameterized by $\mathbf{\Theta}$ on this dataset $\mathcal{D}_{\text{pre}}$. Specifically, we define $\mathrm{TF}_{\mathbf{\Theta}}\left(\cdot\mid\mathcal{H}^{t}_{m}\right)$ as the transformer model that observes the dataset $\mathcal{H}^{t}_{m}$ till round $t$ and then produces a distribution over the actions. Our primary novelty lies in our training procedure which we explain in detail in Section 3.1.
@@ 2.3 Related In-context Learning Algorithms
In this section, we discuss related algorithms for in-context decision-making. For completeness, we describe the DPT and AD training procedure and algorithm now. During training, DPT first samples $m\sim\mathcal{T}_{\text{pre}}$ and then an in-context dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|,m)$. It adds this $\mathcal{H}_{m}$ to the training dataset $\mathcal{H}_{\text{train}}$, and repeats to collect $M_{\text{pre}}$ such training tasks. For each task $m$, DPT  requires the optimal action $a_{m,*}=\mathop{\rm arg\,max}_{a}f(\mathbf{x}(m,a),\bm{\theta}_{m,*})$ where $f(\mathbf{x}(m,a),\bm{\theta}_{m,*})$ is the expected reward for the action $a$ in task $m$. Since the optimal action is usually not known in advance, in Section 4 we introduce a practical variant of DPT  that approximates the optimal action with the best action identified during task interaction. During training DPT minimizes the cross-entropy loss:
@@
where $p({a}_{m,*})\!\in\!\triangle^{\mathcal{A}}$ is a one-hot vector such that $p(j)\!=\!1$ when $j\!=\!{a}_{m,*}$ and 00 otherwise. This loss is then back-propagated and used to update the model parameter $\mathbf{\Theta}$.
@@
where $p({I}_{m,t})$ is a one-hot vector such that $p(j)=1$ when $j={I}_{m,t}$ (the true action taken by the demonstrator) and 00 otherwise. At deployment time, AD selects $I_{t}\sim\mathrm{softmax}^{\tau}_{a}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}^{t}_{m}))$. The objective of AD is to match the performance of the demonstrator. In the next section, we introduce a new method that can improve upon the demonstrator without knowledge of the optimal action.
3 Proposed Algorithm PreDeToR
@@ 3.1 Pre-training Next Reward Prediction
The key idea behind PreDeToR is to leverage the in-context learning ability of transformers to infer the reward of each arm in a given test task. By training this in-context ability on a set of training tasks, the transformer can implicitly learn structure in the task family and exploit this structure to infer rewards without trying every single arm. Thus, in contrast to DPT and AD that output actions directly, PreDeToR outputs a scalar value reward prediction for each arm. To this effect, we append a linear layer of dimension $A$ on top of a causal GPT2 model, denoted by $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$, and use a least-squares loss to train the transformer to predict the reward for each action with these outputs. Note that we use $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ to denote a reward prediction transformer and $\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ as the transformer that predicts a distribution over actions (as in DPT and AD ). At every round $t$ the transformer predicts the next reward for each of the actions $a\in\mathcal{A}$ for the task $m$ based on $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This predicted reward is denoted by $\widehat{r}_{m,t+1}(a)$ for each $a\in\mathcal{A}$.
@@
Exploratory Demonstrator: Observe from the loss definition in (3) that it is calculated from the observed true reward and action from the dataset $\mathcal{H}_{m}$. In order for the transformer to learn accurate reward predictions during training, we require that the weak demonstrator is sufficiently exploratory such that it collects $\mathcal{H}_{m}$ such that $\mathcal{H}_{m}$ contains some reward $r_{m,t}$ for each action $a$. We discuss in detail the impact of the demonstrator on PreDeToR (-$\tau$) training in Section 7.
@@ 3.2 Deploying PreDeToR
At deployment time, PreDeToR learns in-context to predict the mean reward of each arm on an unseen task and acts greedily with respect to this prediction. That is, at deployment time, a new task is sampled, $m\sim\mathcal{T}_{\text{test}}$, and the dataset $\mathcal{H}_{m}^{0}$ is initialized empty. Then at every round $t$, PreDeToR chooses $I_{t}=\mathop{\rm arg\,max}_{a\in\mathcal{A}}\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\widehat{r}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)$ which is the action with the highest predicted reward and $\widehat{r}_{m,t}(a)$ is the predicted reward of action $a$. Note that PreDeToR is a greedy policy and thus may fail to conduct sufficient exploration. To remedy this potential limitation, we also introduce a soft variant, PreDeToR-$\tau$  that chooses $I_{t}\sim\mathrm{softmax}^{\tau}_{a}\left(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\mathbf{\widehat{r}}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)\right)$. For both PreDeToR and PreDeToR-$\tau$, the observed reward $r_{t}(I_{t})$ is added to the dataset $\mathcal{H}_{m}$ and then used to predict the reward at the next round $t+1$. The full pseudocode of using PreDeToR for online interaction is shown in Algorithm 1. In Section A.14, we discuss how PreDeToR (-$\tau$) can be deployed for offline learning.
@@ 4 Empirical Study: Non-Linear Structure
Having introduced PreDeToR, we now investigate its performance in diverse bandit settings compared to other in-context learning algorithms. In our first set of experiments, we use a bandit setting with a common non-linear structure across tasks. Ideally, a good learner would leverage the structure, however, we choose the structure such that no existing algorithms are well-suited to the non-linear structure. This setting is thus a good testbed for establishing that in-context learning can discover and exploit common structure. Moreover, each task only consists of a few rounds of interactions. This setting is quite common in recommender settings where user interaction with the system lasts only for a few rounds and has an underlying non-linear structure (Kwon et al., 2022; Tomkins et al., 2020). We show that PreDeToR achieves lower regret than other in-context algorithms for the non-linear structured bandit setting. We study the performance of PreDeToR in the large horizon setting in Section A.8.
@@
(1) PreDeToR: This is our proposed method shown in Algorithm 1.
(2) PreDeToR-$\tau$: This is the proposed exploratory method shown in Algorithm 1 and we fix $\tau=0.05$.
(3) DPT-greedy: This baseline is the greedy approximation of the DPT algorithm from Lee et al. (2023) which is discussed in Section 2.3. Note that we choose DPT-greedy as a representative example of similar in-context decision-making algorithms studied in Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c; a) all of which require the optimal action (or its greedy approximation). DPT-greedy estimates the optimal arm using the reward estimates for each arm during each task.
(4) AD: This is the Algorithmic Distillation method (Laskin et al., 2022; Lu et al., 2021) discussed in Section 2.3.
@@
We describe in detail the baselines Thomp, LinUCB, and MLinGreedy for interested readers in Section A.2.2.
Outcomes: Before presenting the result we discuss the main outcomes from our experimental results in this section:
@@
Experimental Result: These findings are reported in Figure 1. In Figure 1(a) we show the non-linear bandit setting for horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$. The demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy. Note that for this low data regime (short horizon) the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. The PreDeToR (-$\tau$) has lower regret than LinUCB, and MLinGreedy, which fail to perform well in this non-linear setting due to their algorithmic design and linear feedback assumption. Finally, PreDeToR-$\tau$ performs slightly better than PreDeToR in both settings as it conducts more exploration.
In Figure 1(b) we show the non-linear bandit setting for horizon $n=25$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$ where the norm of the $\bm{\theta}_{m,*}$ determines the reward of the actions which also is a non-linear function $\bm{\theta}_{m,*}$ and action features. This setting is similar to the wheel bandit setting of Riquelme et al. (2018). Again, we observe that PreDeToR has lower cumulative regret than all the other baselines.
Finally in Figure 1(c) and Figure 1(d) we show the performance of PreDeToR against other baselines in real-world datasets Movielens and Yelp. The Movielens dataset consists of more than 32 million ratings of 200,000 users and 80,000 movies (Harper & Konstan, 2015) where each entry consists of user-id, movie-id, rating, and timestamp. The Yelp dataset (Asghar, 2016) consists of ratings of 1300 business categories by 150,000 users. Each entry is summarized as user-id, business-id, rating, and timestamp. Previously structured bandit works (Deshpande & Montanari, 2012; Hong et al., 2023) directly fit a linear structure or low-rank factorization to estimate the $\bm{\theta}_{m,*}$ and simulate the ratings. However, we directly use the user-ids and movie-ids (or business-ids) to build a histogram of ratings per user and calculate the mean rating per movie (or business-id) per task. Define this as the $\{\mu_{m,a}\}_{a=1}^{A}$. This is then used to simulate the rating for $n$ horizon per movie per task where the data collection algorithm is uniform sampling. Note that this does not require estimation of user or movie features, and PreDeToR (-$\tau$) learns to exploit the latent structure of user-movie (or business) rating correlations directly from the data. From Figure 1(c) and Figure 1(d) we see that PreDeToR, and PreDeToR-$\tau$ outperform all the other baselines in these settings.
@@
5 Empirical Study: Linear Structure and Understanding the Exploration of PreDeToR

### v3

@@ Abstract
We study learning to learn for the multi-task structured bandit problem where the goal is to learn a near-optimal algorithm that minimizes cumulative regret. The tasks share a common structure and an algorithm should exploit the shared structure to minimize the cumulative regret for an unseen but related test task. We use a transformer as a decision-making algorithm to learn this shared structure from data collected by a demonstrator on a set of training task instances. Our objective is to devise a training procedure such that the transformer will learn to outperform the demonstrator’s learning algorithm on unseen test task instances. Prior work on pretraining decision transformers either requires privileged information like access to optimal arms or cannot outperform the demonstrator. Going beyond these approaches, we introduce a pre-training approach that trains a transformer network to learn a near-optimal policy in-context. This approach leverages the shared structure across tasks, does not require access to optimal actions, and can outperform the demonstrator. We validate these claims over a wide variety of structured bandit problems to show that our proposed solution is general and can quickly identify expected rewards on unseen test tasks to support effective exploration.
@@ 1 Introduction
    We introduce a new pre-training procedure, PreDeToR, for learning the underlying reward structure and using this to circumvent the issue of requiring access to the optimal (or approximately optimal) action during training time.
@@ 2.1 Preliminaries
Define $n]=\{1,2,\ldots,n\}$. In a multi-task structured bandit setting the expected reward for each action in each task is assumed to be an unknown function of the hidden parameter and action features (Lattimore & Szepesvári, 2020; Gupta et al., 2020). The interaction proceeds iteratively over $n$ rounds for each task $m\in[M]$. At each round $t\in[n]$ for each task $m\in[M]$, the learner selects an action $I_{m,t}\in\mathcal{A}$ and observes the reward $r_{m,t}=f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})+\eta_{m,t}$, where $\bm{\theta}_{m,*}\in\mathbb{R}^{d}$ is the hidden parameter specific to the task $m$ to be learned by the learner. The function $f(\cdot,\cdot)$ is the unknown reward structure. This can be $f(\mathbf{x}(I_{m,t}),\bm{\theta}_{m,*})=\mathbf{x}(I_{m,t})^{\top}\bm{\theta}_{m,*}$ for the linear setting or even more complex correlation between features and $\bm{\theta}_{m,*}$ (Filippi et al., 2010; Abbasi-Yadkori et al., 2011; Riquelme et al., 2018; Lattimore & Szepesvári, 2019; Dong et al., 2021).
@@ 2.2 In-Context Learning Model
Pretraining: Let $\mathcal{T}_{\text{pre}}$ denote the distribution over tasks $m$ at the time of pretraining. Let $\mathcal{D}_{\text{pre}}$ be the distribution over all possible interactions that the $\pi^{w}$ can generate. We first sample a task $m\sim\mathcal{T}_{\text{pre }}$ and then a context $\mathcal{H}_{m}$ which is a sequence of interactions for $n$ rounds conditioned on the task $m$ such that $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|m)$. So $\mathcal{H}_{m}=\{I_{m,t},r_{m,t}\}_{t=1}^{n}$. We call this dataset $\mathcal{H}_{m}$ an in-context dataset as it contains the contextual information about the task $m$. We denote the samples in $\mathcal{H}_{m}$ till round $t$ as $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This dataset $\mathcal{H}_{m}$ can be collected in several ways: (1) random interactions within $m$, (2) demonstrations from an expert, and (3) rollouts of an algorithm. Finally, we train a causal GPT-2 transformer model $\mathrm{TF}$ parameterized by $\mathbf{\Theta}$ on this dataset $\mathcal{D}_{\text{pre}}$. Specifically, we define $\mathrm{TF}_{\mathbf{\Theta}}\left(\cdot\mid\mathcal{H}^{t}_{m}\right)$ as the transformer model that observes the dataset $\mathcal{H}^{t}_{m}$ till round $t$ and then produces a distribution over the actions. Our primary novelty lies in our training procedure which we explain in detail in Section˜3.1.
@@ 2.3 Related In-context Learning Algorithms
In this section, we discuss related algorithms for in-context decision-making. For completeness, we describe the DPT and AD training procedure and algorithm now. During training, DPT first samples $m\sim\mathcal{T}_{\text{pre}}$ and then an in-context dataset $\mathcal{H}_{m}\sim\mathcal{D}_{\text{pre}}(\cdot|,m)$. It adds this $\mathcal{H}_{m}$ to the training dataset $\mathcal{H}_{\text{train}}$, and repeats to collect $M_{\text{pre}}$ such training tasks. For each task $m$, DPT requires the optimal action $a_{m,*}=\mathop{\rm arg\,max}_{a}f(\mathbf{x}(m,a),\bm{\theta}_{m,*})$ where $f(\mathbf{x}(m,a),\bm{\theta}_{m,*})$ is the expected reward for the action $a$ in task $m$. Since the optimal action is usually not known in advance, in Section˜4 we introduce a practical variant of DPT that approximates the optimal action with the best action identified during task interaction. During training DPT minimizes the cross-entropy loss:
@@
where $p({a}_{m,*})\!\in\!\triangle^{\mathcal{A}}$ is a one-hot vector such that p(j)=1p(j)\!=\!1 when $j\!=\!{a}_{m,*}$ and 0 otherwise. This loss is then back-propagated and used to update the model parameter $\mathbf{\Theta}$.
@@
where $p({I}_{m,t})$ is a one-hot vector such that $p(j)=1$ when $j={I}_{m,t}$ (the true action taken by the demonstrator) and 0 otherwise. At deployment time, AD selects $I_{t}\sim\mathrm{softmax}^{\tau}_{a}(\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}^{t}_{m}))$. The objective of AD is to match the performance of the demonstrator. In the next section, we introduce a new method that can improve upon the demonstrator without knowledge of the optimal action.
3 The PreDeToRAlgorithm
@@ 3.1 Pre-training Next Reward Prediction
The key idea behind PreDeToR is to leverage the in-context learning ability of transformers to infer the reward of each arm in a given test task. By training this in-context ability on a set of training tasks, the transformer can implicitly learn structure in the task family and exploit this structure to infer rewards without trying every single arm. Hence PreDeToR requires access to all the finite set of arms. Note that the AD, and DPT only require access to the arms selected by the demonstrator. In contrast to DPT and AD that output actions directly, PreDeToR outputs a scalar value reward prediction for each arm. To this effect, we append a linear layer of dimension $A$ on top of a causal GPT2 model, denoted by $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$, and use a least-squares loss to train the transformer to predict the reward for each action with these outputs. Note that we use $\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ to denote a reward prediction transformer and $\mathrm{TF}_{\mathbf{\Theta}}(\cdot|\mathcal{H}_{m})$ as the transformer that predicts a distribution over actions (as in DPT and AD ). At every round $t$ the transformer predicts the next reward for each of the actions $a\in\mathcal{A}$ for the task $m$ based on $\mathcal{H}_{m}^{t}=\{I_{m,s},r_{m,s}\}_{s=1}^{t-1}$. This predicted reward is denoted by $\widehat{r}_{m,t+1}(a)$ for each $a\in\mathcal{A}$.
@@
Exploratory Demonstrator: Observe from the loss definition in (3) that it is calculated from the observed true reward and action from the dataset $\mathcal{H}_{m}$. In order for the transformer to learn accurate reward predictions during training, we require that the weak demonstrator is sufficiently exploratory such that it collects $\mathcal{H}_{m}$ such that $\mathcal{H}_{m}$ contains some reward $r_{m,t}$ for each action $a$. We discuss in detail the impact of the demonstrator on PreDeToR (-$\tau$) training in Section˜7.
@@ 3.2 Deploying PreDeToR
At deployment time, PreDeToR learns in-context to predict the mean reward of each arm on an unseen task and acts greedily with respect to this prediction. That is, at deployment time, a new task is sampled, $m\sim\mathcal{T}_{\text{test}}$, and the dataset $\mathcal{H}_{m}^{0}$ is initialized empty. Then at every round $t$, PreDeToR chooses $I_{t}=\mathop{\rm arg\,max}_{a\in\mathcal{A}}\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\widehat{r}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)$ which is the action with the highest predicted reward and $\widehat{r}_{m,t}(a)$ is the predicted reward of action $a$. Note that PreDeToR is a greedy policy and thus may fail to conduct sufficient exploration. To remedy this potential limitation, we also introduce a soft variant, PreDeToR-$\tau$  that chooses $I_{t}\sim\mathrm{softmax}^{\tau}_{a}\left(\mathrm{TF^{\mathbf{r}}}_{\mathbf{\Theta}}\left(\mathbf{\widehat{r}}_{m,t}(a)\mid\mathcal{H}^{t}_{m}\right)\right)$. For both PreDeToR and PreDeToR-$\tau$, the observed reward $r_{t}(I_{t})$ is added to the dataset $\mathcal{H}_{m}$ and then used to predict the reward at the next round $t+1$. The full pseudocode of using PreDeToR for online interaction is shown in Algorithm˜1. In Section˜A.14, we discuss how PreDeToR (-$\tau$) can be deployed for offline learning. We also highlight that PreDeToR needs to forward-pass $|A|$ times to select the best arm during evaluation, and hence suffers from more computational overhead with a large action space, compared to AD or DPT.
@@ 4 Empirical Study: Non-Linear Structure
Having introduced PreDeToR, we now investigate its performance in diverse bandit settings compared to other in-context learning algorithms. In our first set of experiments, we use a bandit setting with a common non-linear structure across tasks. Ideally, a good learner would leverage the structure, however, we choose the structure such that no existing algorithms are well-suited to the non-linear structure. This setting is thus a good testbed for establishing that in-context learning can discover and exploit common structure. Moreover, each task only consists of a few rounds of interactions. This setting is quite common in recommender settings where user interaction with the system lasts only for a few rounds and has an underlying non-linear structure (Kwon et al., 2022; Tomkins et al., 2020). We show that PreDeToR achieves lower regret than other in-context algorithms for the non-linear structured bandit setting. We study the performance of PreDeToR in the large horizon setting in Section˜A.8.
@@
(1) PreDeToR: This is our proposed method shown in Algorithm˜1.
(2) PreDeToR-$\tau$: This is the proposed exploratory method shown in Algorithm˜1 and we fix $\tau=0.05$.
(3) DPT-greedy: This baseline is the greedy approximation of the DPT algorithm from Lee et al. (2023) which is discussed in Section˜2.3. Note that we choose DPT-greedy as a representative example of similar in-context decision-making algorithms studied in Lee et al. (2023); Sinii et al. (2023); Lin et al. (2023); Ma et al. (2023); Liu et al. (2023c; a) all of which require the optimal action (or its greedy approximation). DPT-greedy estimates the optimal arm using the reward estimates for each arm during each task.
(4) AD: This is the Algorithmic Distillation method (Laskin et al., 2022; Lu et al., 2021) discussed in Section˜2.3.
@@
We describe in detail the baselines Thomp, LinUCB, and MLinGreedy for interested readers in Section˜A.2.2.
Outcomes: First, we discuss the main outcomes from our experimental results in this section:
@@
Experimental Result: These findings are reported in Figure˜1. In Figure˜1(a) we show the non-linear bandit setting for horizon $n=50$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$. The demonstrator $\pi^{w}$ is the Thomp algorithm. We observe that PreDeToR (-$\tau$) has lower cumulative regret than DPT-greedy. Note that for this low data regime (short horizon) the DPT-greedy does not have a good estimation of $\widehat{a}_{m,*}$ which results in a poor prediction of optimal action $\widehat{a}_{m,t,*}$. This results in higher regret. The PreDeToR (-$\tau$) has lower regret than LinUCB, and MLinGreedy, which fail to perform well in this non-linear setting due to their algorithmic design and linear feedback assumption. Finally, PreDeToR-$\tau$ performs slightly better than PreDeToR in both settings as it conducts more exploration.
In Figure˜1(b) we show the non-linear bandit setting for horizon $n=25$, $M_{\text{pre}}=100000$, $M_{\text{test}}=200$, $A=6$, and $d=2$ where the norm of the $\bm{\theta}_{m,*}$ determines the reward of the actions which also is a non-linear function $\bm{\theta}_{m,*}$ and action features. This setting is similar to the wheel bandit setting of Riquelme et al. (2018). Again, we observe that PreDeToR has lower cumulative regret than all the other baselines.
Finally in Figure˜1(c) and Figure˜1(d) we show the performance of PreDeToR against other baselines in real-world datasets Movielens and Yelp. The Movielens dataset consists of more than 32 million ratings of 200,000 users and 80,000 movies (Harper & Konstan, 2015) where each entry consists of user-id, movie-id, rating, and timestamp. The Yelp dataset (Asghar, 2016) consists of ratings of 1300 business categories by 150,000 users. Each entry is summarized as user-id, business-id, rating, and timestamp. Previously structured bandit works (Deshpande & Montanari, 2012; Hong et al., 2023) directly fit a linear structure or low-rank factorization to estimate the $\bm{\theta}_{m,*}$ and simulate the ratings. However, we directly use the user-ids and movie-ids (or business-ids) to build a histogram of ratings per user and calculate the mean rating per movie (or business-id) per task. Define this as the $\{\mu_{m,a}\}_{a=1}^{A}$. This is then used to simulate the rating for $n$ horizon per movie per task where the data collection algorithm is uniform sampling. Note that this does not require estimation of user or movie features, and PreDeToR (-$\tau$) learns to exploit the latent structure of user-movie (or business) rating correlations directly from the data. From Figure˜1(c) and Figure˜1(d) we see that PreDeToR, and PreDeToR-$\tau$ outperform all the other baselines in these settings. In the next section, we study the simplified linear setting to show that PreDeToR is indeed exploiting the latent structure to minimize the cumulative regret.
@@
5 Empirical Study: Linear Structure and Understanding PreDeToR’s Exploration